24. Find the parametric and symmetric equations of the line that passes through the point (2, 1,-3) and is perpendicular to the planex+2y+2- 13. 14 marks] 25. Find the following quantities: 14 marks] a. (1+2)8- 5-5 b 23 (52 -1+i
Answer #1

25) a.

conjugate(1 +2i)(3 - i)

(1 +2i)(3 - i) = 3 +1 + 6i +2 = 6+6i

conjugate (6+6i) = 6 - 6i

conjugate(1 +2i)(3 - i) = 6 -6i

b. (2 +3i)/( -1 +i)

Multiply num. and den. with conjugate of denominator i.e -1 - i

So, ( 2 +3i)(-1 - i)/(1 +1)

(-2 -2i -3i +3)/2

(1 -5i)/2

1/2 - 5i/2